# A globalized inexact semismooth Newton method for strongly convex optimal control problems

### v3

@@ 6.2 Examples
Clearly, this function is piecewise affine linear, and 2 is satisfied. For this particular choice of $\tilde{g}$, we have $\tilde{g}_{\textrm{box}}/\alpha=\tilde{g}_{\textrm{box}}$.
@@
This is again a piecewise affine linear map, so that 2 is satisfied. The proximal operator of $\tilde{g}_{L^{1}}/\alpha$ is obtained by replacing $\beta$ with $\beta/\alpha$ in the formula for $\operatorname{prox}_{\tilde{g}_{L^{1}}}$.
@@
and hence, 2 is satisfied. Again, the proximal operator of $(\tilde{g}_{L^{1}}+\tilde{g}_{\textrm{box}})/\alpha$ is obtained by replacing $\beta$ with $\beta/\alpha$ in the formula above.
@@
Then $T^{\prime}=\operatorname{prox}_{g}$ and $T^{\prime\prime}=\partial\operatorname{prox}_{g}$ satisfy condition (3.2) by [23, Proposition 4.1]. In addition, $\operatorname{prox}_{g}$ is Bouligand differentiable due to [23, Proposition 3.4].
@@
This operator is semismooth from $L^{p}(\Omega;L^{2}(I))$ to $L^{2}(\Omega;L^{2}(I))$ for all $p>2$, [27, Lemma 3.22], [13, Lemma 3.2]. In addition, it is easy to check that $\operatorname{prox}_{g/\alpha}$ is Bouligand differentiable using the chain rule for the Bouligand derivative [32, Corollary A.4].
@@
Additional examples that satisfy these assumptions are multi-bang problems [4], where $g$ is given as the convex hull of
@@
In addition, also vector-valued controls can be considered [5]. Since $u=\operatorname{prox}_{g/\alpha}(v)$ if and only if $0\in\alpha(u-v)+\partial g(u)$, implicit function theorems [21] can be used to prove the semismoothness of proximal maps.
@@ 7 Variational discretization
For computations, one has to choose finite-dimensional approximations of the infinite-dimensional problem (1.1). The application of the semismooth Newton method to discretized problems is straight-forward. In this section, we will focus on the variational discretization of [15]. We will consider the following optimal control problem: Minimize
@@
where $\Omega\subseteq\mathbb{R}^{d}$ is a bounded domain, $z\in L^{2}(\Omega)$ is given, and $\tilde{g}$$\mathbb{R}\to\bar{\mathbb{R}}$ is as in Section 6. The state space is discretized using piecewise linear and continuous finite element functions: let $V_{h}\subseteq H^{1}_{0}(\Omega)$ be the corresponding finite element subspace. Given $u\in L^{2}(\Omega)$, let $S_{h}u:=y_{h}\in V_{h}$ denote the solution of
$\int_{\Omega}\nabla y_{h}\cdot\nabla v_{h}\,\text{{}{}d}x=\int_{\Omega}uv_{h}\,\text{{}{}d}x\quad\forall v_{h}\in V_{h},$ (7.1)
@@
$\xi-z+S_{h}\operatorname{prox}_{g/\alpha}(S_{h}^{*}\xi/\alpha)=0.$ (7.2)
The crucial observation of [15] is the following: in this equation $S_{h}$ is applied to elements of the type $\operatorname{prox}_{g/\alpha}(q_{h})$, where $q_{h}$ is a piecewise linear finite element function. That is, if $\operatorname{prox}_{g/\alpha}$ has a favorable structure, then $S_{h}\operatorname{prox}_{g/\alpha}(q_{h})$ can be evaluated numerically even if $\operatorname{prox}_{g/\alpha}(q_{h})$ does not belong to a finite element subspace. This is the case for the first four examples discussed in Section 6.2. For instance, if $g$ models box constraints on the control, then $\operatorname{prox}_{g/\alpha}(q_{h})$ is piecewise linear. This piecewise linear function has kinks in the interior of elements in general, nevertheless the right-hand side for $u=\operatorname{prox}_{g/\alpha}(q_{h})$ in (7.1) can be computed exactly [15, 16].
In addition, when applying the semismooth Newton method to the dual equation (7.2), these kinks do not accumulate. This is not the case when applying the semismooth Newton to the primal equation as in Remark 5.15. To overcome this difficulty, [17] suggested to apply the semismooth Newton method to the normal map equation as in Remark 5.16. The variational discretization is not limited to piecewise linear finite elements, in [36] the application to finite elements of polynomial degree 2 is discussed.
@@ 8 Numerical experiments
$J(y,u):=\frac{1}{2}\|y-z\|_{L^{2}(\Omega)}^{2}+\frac{\alpha}{2}\|u\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}\tilde{g}(u(x))\,\text{{}{}d}x$ (8.1)
@@
$\int_{\Omega}\nabla y\cdot\nabla v\,\text{{}{}d}x=\int_{\Omega}uv\,\text{{}{}d}x\quad\forall v\in H^{1}_{0}(\Omega),$ (8.2)
where $\Omega=(0,1)^{2}$, $z\in L^{2}(\Omega)$ is given, and $\tilde{g}$$\mathbb{R}\to\bar{\mathbb{R}}$ is as in Section 6. Here, (8.2) is the weak formulation of $\Delta y=u$ with homogeneous Dirichlet boundary conditions.
@@
8.1 Example 1: box constraints and $L^{1}$-cost
@@
The parameter $\alpha$ is set to $\alpha=10^{-5}$ if not mentioned otherwise. The (local) semismooth Newton method for this problem was analyzed in [37]. As argued in Section 6.2, $g$ satisfies 2. Due to the smoothing properties of $S^{*}$, the assumptions of Lemma 6.5 are satisfied, so that 1 holds for this example. As desired state we used
@@
In the implementation, we used the following values for the parameters of Algorithm 2:
@@
which is slightly stronger than the one which is used in the convergence analysis, and is inspired by the choices taken in [25]. In addition, we stopped the iteration if the threshold in the linesearch is smaller than the machine epsilon, i.e.,
$\left|\langle d_{k},\nabla\Phi(\xi_{k})\rangle\right|\leq\tt{eps(dual)},$ (8.3)
@@
where the first inequality is due to convexity. Hence, $\Phi(\xi_{k}+t_{k}d_{k})$ lies in an interval of length $t_{k}\left|\langle d_{k},\nabla\Phi(\xi_{k})\rangle\right|$. If the condition (8.3) is satisfied, then the only floating point number in this interval is $\Phi(\xi_{k})$, hence it is impossible to decrease the dual functional value any further. In addition, Corollary 5.6 implies that $\|\nabla\Phi(\xi_{k})\|_{Y}$ is small in this situation.
The optimal control $u=\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$ computed on a mesh consisting of $20000$ triangles and with mesh-size $h=0.014$ is depicted in Figure 1. The right-hand plot in Figure 1 shows $\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi_{k}/\alpha)$, where $\xi_{k}$ is the final iterate of the dual variable. In this plot, dark regions correspond to $\partial\operatorname{prox}_{g/\alpha}=0$, which means that the control is either zero or at the bounds $\pm R$, while light regions correspond to $\partial\operatorname{prox}_{g/\alpha}=0$, which are regions, where the control is between zero and the bounds.
@@
The results for computations on different meshes can be found in Table 1. The mesh size is denoted by $h$. The finest mesh consists of $2\cdot 10^{6}$ triangles. On all meshes the globalized semismooth Newton method needed 8 iterations. The final value of the dual objective $\Phi$ together with the value of the final residual and primal-dual gap can be found in Table 1 as well. In all cases the iteration stopped because the condition (8.3) was satisfied. Since the primal-dual gap is small compared to $\Phi$, the optimal primal objective value equals $\Phi$. In addition, the number of iterations of the globalized semismooth Newton method (it) as well as the total number of iterations of the inner conjugate gradient method (cg) almost stay constant for the different discretizations.
@@
We also performed experiments with varying parameter $\alpha$ on a fixed discretization with $h=2.83\cdot 10^{-3}$. The results can be found in Table 2. As one expects, the iteration number increases with decreasing $\alpha$. The third column of Table 2 reports the value $\|\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)\|_{L^{1}(\Omega)}$, which corresponds to the size of the inactive set, i.e., the size of the set, where $\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$ takes values not in $\{-R,0,+R\}$.
@@
In addition, we conducted experiments using the variational discretization of [15] as described in Section 7. We studied again varying $\alpha$ on a fixed mesh with $h=2.83\cdot 10^{-3}$. The results can be found in Table 3. While the results are comparable to those in Table 2 it is surprising that the iteration numbers grow only moderately with decreasing $\alpha$. So it seems that the variational discretization should be favored to the piecewise constant discretization at least for small values of $\alpha$. This might be due to the fact that the variational discretization can better resolve the boundary of active sets.
@@ 8.2 Example 2: singular control problem
Hence, the control is used to counter the unwanted perturbation $f$ in the equation. This example is adapted from an example in [35, Section 4.4]. It is challenging to solve for small values of $\alpha$, as for $\alpha\searrow 0$ the corresponding solutions are not of bang-bang type.
@@
We used the same discretization and setup as in the previous experiment. The solution for two different values of $\alpha\in\{10^{-8},10^{-10}\}$ and for a discretization with $h=2.83\cdot 10^{-3}$ can be found in Figure 2. The control is zero on a large part of the domain. On the interface of the regions between $u=0$ and $u\neq 0$ the control exhibits scattering behavior: it oscillates between upper and lower bounds, where the oscillations increase with decreasing mesh size and parameter $\alpha$. The results of computations for different values of $\alpha$ are reported in Table 4. As one can see, the iteration number grows much stronger with decreasing $\alpha$ than in the previous experiment. Most of these iterations are spent in damping oscillations of the control variable in the region, where the control eventually is zero. In addition, the inner CG iteration also took considerably more steps than for the previous experiment.
@@
That lead us to consider a nested iteration approach: here the iteration for some value $\alpha=10^{-(k+1)}$ is started with the result for the value $\alpha=10^{-k}$. The results can be found in Table 5. There the column ’iterations’ denotes the number of iterations when started with the solution of the previous iterations. For this warm-started semismooth Newton the number of iterations per value of $\alpha$ stays constant. In addition, the cumulated iteration numbers are still less than those for the experiment in Table 4.

### v4

@@ 6.2 Examples
Clearly, this function is piecewise affine linear, and ˜2 is satisfied. For this particular choice of $\tilde{g}$, we have $\tilde{g}_{\textrm{box}}/\alpha=\tilde{g}_{\textrm{box}}$.
@@
This is again a piecewise affine linear map, so that ˜2 is satisfied. The proximal operator of $\tilde{g}_{L^{1}}/\alpha$ is obtained by replacing $\beta$ with $\beta/\alpha$ in the formula for $\operatorname{prox}_{\tilde{g}_{L^{1}}}$.
@@
and hence, ˜2 is satisfied. Again, the proximal operator of $(\tilde{g}_{L^{1}}+\tilde{g}_{\textrm{box}})/\alpha$ is obtained by replacing $\beta$ with $\beta/\alpha$ in the formula above.
@@
Then $T^{\prime}=\operatorname{prox}_{g}$ and $T^{\prime\prime}=\partial\operatorname{prox}_{g}$ satisfy condition (3.2) by [KunischWachsmuth2013, Proposition 4.1]. In addition, $\operatorname{prox}_{g}$ is Bouligand differentiable due to [KunischWachsmuth2013, Proposition 3.4].
@@
This operator is semismooth from $L^{p}(\Omega;L^{2}(I))$ to $L^{2}(\Omega;L^{2}(I))$ for all $p>2$, [PieperDiss, Lemma 3.22], [HerzogStadlerWachsmuth2012, Lemma 3.2]. In addition, it is easy to check that $\operatorname{prox}_{g/\alpha}$ is Bouligand differentiable using the chain rule for the Bouligand derivative [Robinson1987, Corollary A.4].
@@
Additional examples that satisfy these assumptions are multi-bang problems [ClasonKunisch2014], where $g$ is given as the convex hull of
@@
In addition, also vector-valued controls can be considered [ClasonTamelingWirth2021]. Since $u=\operatorname{prox}_{g/\alpha}(v)$ if and only if $0\in\alpha(u-v)+\partial g(u)$, implicit function theorems [Kruse2018] can be used to prove the semismoothness of proximal maps.
@@ 7 Variational discretization
For computations, one has to choose finite-dimensional approximations of the infinite-dimensional problem (1.1). The application of the semismooth Newton method to discretized problems is straight-forward. In this section, we will focus on the variational discretization of [Hinze2005]. We will consider the following optimal control problem: Minimize
@@
where $\Omega\subseteq\mathbb{R}^{d}$ is a bounded domain, $z\in L^{2}(\Omega)$ is given, and $\tilde{g}$$\mathbb{R}\to\bar{\mathbb{R}}$ is as in Section˜6. The state space is discretized using piecewise linear and continuous finite element functions: let $V_{h}\subseteq H^{1}_{0}(\Omega)$ be the corresponding finite element subspace. Given $u\in L^{2}(\Omega)$, let $S_{h}u:=y_{h}\in V_{h}$ denote the solution of
(7.1) $\int_{\Omega}\nabla y_{h}\cdot\nabla v_{h}\,\text{{}{}d}x=\int_{\Omega}uv_{h}\,\text{{}{}d}x\quad\forall v_{h}\in V_{h},$
@@
(7.2) $\xi-z+S_{h}\operatorname{prox}_{g/\alpha}(S_{h}^{*}\xi/\alpha)=0.$
The crucial observation of [Hinze2005] is the following: in this equation $S_{h}$ is applied to elements of the type $\operatorname{prox}_{g/\alpha}(q_{h})$, where $q_{h}$ is a piecewise linear finite element function. That is, if $\operatorname{prox}_{g/\alpha}$ has a favorable structure, then $S_{h}\operatorname{prox}_{g/\alpha}(q_{h})$ can be evaluated numerically even if $\operatorname{prox}_{g/\alpha}(q_{h})$ does not belong to a finite element subspace. This is the case for the first four examples discussed in Section˜6.2. For instance, if $g$ models box constraints on the control, then $\operatorname{prox}_{g/\alpha}(q_{h})$ is piecewise linear. This piecewise linear function has kinks in the interior of elements in general, nevertheless the right-hand side for $u=\operatorname{prox}_{g/\alpha}(q_{h})$ in (7.1) can be computed exactly [Hinze2005, HinzePinnauUlbrichUlbrich2009].
In addition, when applying the semismooth Newton method to the dual equation (7.2), these kinks do not accumulate. This is not the case when applying the semismooth Newton to the primal equation as in Remark˜5.26. To overcome this difficulty, [HinzeVierling_2012spp] suggested to apply the semismooth Newton method to the normal map equation as in Remark˜5.27. The variational discretization is not limited to piecewise linear finite elements, in [SevillaWachsmuth2010] the application to finite elements of polynomial degree 2 is discussed.
@@ 8 Numerical experiments
(8.1) $J(y,u):=\frac{1}{2}\|y-z\|_{L^{2}(\Omega)}^{2}+\frac{\alpha}{2}\|u\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}\tilde{g}(u(x))\,\text{{}{}d}x$
@@
(8.2) $\int_{\Omega}\nabla y\cdot\nabla v\,\text{{}{}d}x=\int_{\Omega}uv\,\text{{}{}d}x\quad\forall v\in H^{1}_{0}(\Omega),$
where $\Omega=(0,1)^{2}$, $z\in L^{2}(\Omega)$ is given, and $\tilde{g}$$\mathbb{R}\to\bar{\mathbb{R}}$ is as in Section˜6. Here, (8.2) is the weak formulation of $\Delta y=u$ with homogeneous Dirichlet boundary conditions.
@@
8.1 Example 1: box constraints and $\scriptstyle L^{1}$-cost
@@
The parameter $\alpha$ is set to $\alpha=10^{-5}$ if not mentioned otherwise. The (local) semismooth Newton method for this problem was analyzed in [Stadler2009]. As argued in Section˜6.2, $g$ satisfies ˜2. Due to the smoothing properties of $S^{*}$, the assumptions of Lemma˜6.8 are satisfied, so that ˜1 holds for this example. As desired state we used
@@
In the implementation, we used the following values for the parameters of Algorithm˜2:
@@
which is slightly stronger than the one which is used in the convergence analysis, and is inspired by the choices taken in [MilzarekSchaippUlbrich2024]. In addition, we stopped the iteration if the threshold in the linesearch is smaller than the machine epsilon, i.e.,
(8.3) $\left|\langle d_{k},\nabla\Phi(\xi_{k})\rangle\right|\leq\texttt{eps(dual)},$
@@
where the first inequality is due to convexity. Hence, $\Phi(\xi_{k}+t_{k}d_{k})$ lies in an interval of length $t_{k}\left|\langle d_{k},\nabla\Phi(\xi_{k})\rangle\right|$. If the condition (8.3) is satisfied, then the only floating point number in this interval is $\Phi(\xi_{k})$, hence it is impossible to decrease the dual functional value any further. In addition, Corollary˜5.9 implies that $\|\nabla\Phi(\xi_{k})\|_{Y}$ is small in this situation.
The optimal control $u=\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$ computed on a mesh consisting of $20000$ triangles and with mesh-size $h=0.014$ is depicted in Figure˜1. The right-hand plot in Figure˜1 shows $\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi_{k}/\alpha)$, where $\xi_{k}$ is the final iterate of the dual variable. In this plot, dark regions correspond to $\partial\operatorname{prox}_{g/\alpha}=0$, which means that the control is either zero or at the bounds $\pm R$, while light regions correspond to $\partial\operatorname{prox}_{g/\alpha}=0$, which are regions, where the control is between zero and the bounds.
@@
The results for computations on different meshes can be found in Table˜1. The mesh size is denoted by $h$. The finest mesh consists of $2\cdot 10^{6}$ triangles. On all meshes the globalized semismooth Newton method needed 8 iterations. The final value of the dual objective $\Phi$ together with the value of the final residual and primal-dual gap can be found in Table˜1 as well. In all cases the iteration stopped because the condition (8.3) was satisfied. Since the primal-dual gap is small compared to $\Phi$, the optimal primal objective value equals $\Phi$. In addition, the number of iterations of the globalized semismooth Newton method (it) as well as the total number of iterations of the inner conjugate gradient method (cg) almost stay constant for the different discretizations.
@@
We also performed experiments with varying parameter $\alpha$ on a fixed discretization with $h=2.83\cdot 10^{-3}$. The results can be found in Table˜2. As one expects, the iteration number increases with decreasing $\alpha$. The third column of Table˜2 reports the value $\|\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)\|_{L^{1}(\Omega)}$, which corresponds to the size of the inactive set, i.e., the size of the set, where $\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$ takes values not in $\{-R,0,+R\}$.
@@
In addition, we conducted experiments using the variational discretization of [Hinze2005] as described in Section˜7. We studied again varying $\alpha$ on a fixed mesh with $h=2.83\cdot 10^{-3}$. The results can be found in Table˜3. While the results are comparable to those in Table˜2 it is surprising that the iteration numbers grow only moderately with decreasing $\alpha$. So it seems that the variational discretization should be favored to the piecewise constant discretization at least for small values of $\alpha$. This might be due to the fact that the variational discretization can better resolve the boundary of active sets.
@@ 8.2 Example 2: singular control problem
Hence, the control is used to counter the unwanted perturbation $f$ in the equation. This example is adapted from an example in [Rotin, Section 4.4]. It is challenging to solve for small values of $\alpha$, as for $\alpha\searrow 0$ the corresponding solutions are not of bang-bang type.
@@
We used the same discretization and setup as in the previous experiment. The solution for two different values of $\alpha\in\{10^{-8},10^{-10}\}$ and for a discretization with $h=2.83\cdot 10^{-3}$ can be found in Figure˜2. The control is zero on a large part of the domain. On the interface of the regions between $u=0$ and $u\neq 0$ the control exhibits scattering behavior: it oscillates between upper and lower bounds, where the oscillations increase with decreasing mesh size and parameter $\alpha$. The results of computations for different values of $\alpha$ are reported in Table˜4. As one can see, the iteration number grows much stronger with decreasing $\alpha$ than in the previous experiment. Most of these iterations are spent in damping oscillations of the control variable in the region, where the control eventually is zero. In addition, the inner CG iteration also took considerably more steps than for the previous experiment.
@@
That lead us to consider a nested iteration approach: here the iteration for some value $\alpha=10^{-(k+1)}$ is started with the result for the value $\alpha=10^{-k}$. The results can be found in Table˜5. There the column ’iterations’ denotes the number of iterations when started with the solution of the previous iterations. For this warm-started semismooth Newton the number of iterations per value of $\alpha$ stays constant. In addition, the cumulated iteration numbers are still less than those for the experiment in Table˜4, and the inner CG iteration numbers do not grow as severely as in that setting.